[Brahms-l] Shift report 20040511 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue May 11 2004 - 23:32:06 EDT
------------------------------------------------------------
Date: Tue May 11 23:32:06 2004, Shift: 16:00-24:00
Supervisor: Michael, with: 
Log book#15 pages 118 - 118

Shift Summary:
------------------------------------------------------------
1) Beam on for 2 hours or so until fire alarm in Phenix and then control problems.
2) No data only pedestal problems
3) All magnets tripped off shortly after putting them on. This took most of the evening to fix.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 11 23:32:26 2004

This archive was generated by hypermail 2.1.8 : Tue May 11 2004 - 23:32:49 EDT